OpenStudy (anonymous):

The graph of y ≤ -x2 - 6x - 5 contains which of the following points

ganeshie8 (ganeshie8):

where are the `following points` ?

OpenStudy (anonymous):

A. (-3, 5) B. (-3, 4) C. (1, 6) D. (2, 0)

OpenStudy (anonymous):

@jojo9999

OpenStudy (cwrw238):

jsut plug in the x and y values of each point and see if they fit the inequality eg the point (-3,5):- x = -3 and y = 5 replace x and y by these values ; y <= -x2 - 6x - 5 5 <= -(-3)^2 - 6(-3) - 5 5 <= -9 + 18 - 5 = 4 this isn't true so (-3,5) is not contained by the graph